A food manufacturing company produces oatmeal cookies that have a sugar content that



is approximately normally distributed. The mean sugar content is 1.1 grams with a standard



deviation of 0.15 gram. Determine the probability that a random sample of 10 oatmeal cookies



will have an average sugar content of greater than 1.2 grams.

1
Expert's answer
2022-03-31T07:47:19-0400

We have a normal distribution, "\mu=1.1, \sigma=0.15, n=10."

Let's convert it to the standard normal distribution,

"z=\cfrac{\bar{x}-\mu}{\sigma/\sqrt{n}}=\cfrac{1.2-1.1}{0.15/\sqrt{10}}=2.11,"

"P(\bar{X}>1.2)=1-P(\bar{X}<1.2)=\\
=1-P(Z<2.11)="

"=1-0.9826=0.0174" ​(from z-table).